# Learning Theory for Kernel Bilevel Optimization

### v1

@@ Appendix B Gradient Estimators
Let $\omega\in\mathbb{R}^{d}$. Recall the expression of ^$\widehat{\mathcal{F}}(\omega)$:
^$\widehat{\mathcal{F}}(\omega)\coloneqq\frac{1}{m}\sum_{j=1}^{m}\ell_{out}{\left(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j}\right)}$
@@
$\hat{h}_{\omega}=\sum_{i=1}^{n}(\hat{\gamma}_{\omega})_{i}K(x_{i},\cdot).$ (10)
Hence, finding $\hat{h}_{\omega}$ amounts to minimizing $\widehat{L}_{in}(\omega,h)$ over the span of $(K(x_{1},\cdot),\ldots,K(x_{n},\cdot))$, i.e., over functions $h^{\gamma}$ of the form $h^{\gamma}=\sum_{i=1}^{n}\gamma_{i}K(x_{i},\cdot)$ for $\gamma\in\mathbb{R}^{n}$. Restricting the objective to such functions results in the following inner optimization problem which is finite-dimensional:
$\hat{\gamma}_{\omega}\coloneqq\operatorname*{arg\,min}_{\gamma\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell_{in}{\left(\omega,(\operatorname{\mathbf{K}}\gamma)_{i},y_{i}\right)}+\frac{\lambda}{2}\gamma^{\top}\operatorname{\mathbf{K}}\gamma,$
where we used that $(h^{\gamma}(x_{i}))_{1\leq i\leq n}=\operatorname{\mathbf{K}}\gamma$ and ${\left\lVert h^{\gamma}\right\rVert}^{2}_{\mathcal{H}}=\gamma^{\top}\operatorname{\mathbf{K}}\gamma$. Similarly, using that $(h^{\gamma}(\tilde{x}_{j}))_{1\leq j\leq m}=\operatorname{\mathbf{\overline{K}}}\gamma$, we can express ^$\widehat{\mathcal{F}}(\omega)$ as follows:
^$\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\sum_{j=1}^{m}\ell_{out}{\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\gamma}_{\omega})_{j},\tilde{y}_{j}\right)}.$
@@
^$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}{\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\gamma}_{\omega})_{j},\tilde{y}_{j}\right)}+\frac{1}{m}\sum_{j=1}^{m}(\partial_{\omega}\hat{\gamma}_{\omega}\operatorname{\mathbf{\overline{K}}}^{\top})_{j}\partial_{v}\ell_{out}{\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\gamma}_{\omega})_{j},\tilde{y}_{j}\right)},$
where $\partial_{\omega}\hat{\gamma}_{\omega}$ denotes the Jacobian of $\hat{\gamma}_{\omega}$. We can further express the above equation in matrix form to get:
^$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{m}\partial_{\omega}\hat{\gamma}_{\omega}\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}}.$ (11)
Moreover, an application of the implicit function theorem [^1] In the case where the matrix $\operatorname{\mathbf{K}}$ is non-invertible, one needs to restrict $\gamma$ to the orthogonal complement of the null space of $\operatorname{\mathbf{K}}$. Such a restriction is valid since the resulting solution $\hat{h}_{\omega}$ will not depend on the component belonging to the null space of $\operatorname{\mathbf{K}}$. allows to directly express the Jacobian $\partial_{\omega}\hat{\gamma}_{\omega}$ as a solution of the following linear system obtained by differentiating the optimality condition for $\hat{\gamma}_{\omega}$ w.r.t. $\omega$:
@@
A solution of the form $\partial_{\omega}\hat{\gamma}_{\omega}=-\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}$ always exists by invertibility of the matrix $\operatorname{\mathbf{M}}$. The result follows after replacing $\partial_{\omega}\hat{\gamma}_{\omega}$ by $-\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}$ in Equation 11. ∎
Lemma B.2 (Estimator of the total functional gradient).
@@
^$\widehat{\nabla\mathcal{F}}(\omega)=$ $\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}).$
Then, under Assumptions (B), (D), (A) and (C), ^$\widehat{\nabla\mathcal{F}}(\omega)$ admits the following expression:
^$\widehat{\nabla\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{n}\operatorname{\mathbf{D_{\omega,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}\in\mathbb{R}^{d},$
where $\operatorname{\mathbf{D_{v}^{out}}},\operatorname{\mathbf{D_{v,v}^{in}}},\operatorname{\mathbf{D_{\omega}^{out}}},\operatorname{\mathbf{D_{\omega,v}^{in}}}$ are the same matrices given in Proposition B.1, while, $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\in\mathbb{R}^{n}$ and $\hat{\beta}_{\omega}\in\mathbb{R}$ are solutions to the linear system:
@@
where the vector $\mathbf{u}$ and matrix $\operatorname{\mathbf{M}}$ are the same as in Proposition B.1, while $p$ and $v$ are non-negative scalars.
@@
Let $\omega\in\mathbb{R}^{d}$. We start by providing an expression of $\hat{a}_{\omega}$ as a linear combination of the kernel evaluated at the inner training points $x_{i}$, i.e., $K(x_{i},\cdot)$, and some element $\xi\in\mathcal{H}$ that we will characterize shortly. From it, we will obtain the expression of ^$\widehat{\nabla\mathcal{F}}(\omega)$.
Expression of $\hat{a}_{\omega}$. Recall that $\hat{a}_{\omega}$ is the unique minimizer of $\widehat{L}_{adj}$ in Equation 5, which admits, for any $a\in\mathcal{H}$, the following simple expression by the reproducing property:
$\widehat{L}_{adj}(\omega,a)=\frac{1}{2n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})a^{2}(x_{i})+\frac{1}{m}\left\langle a,\overbrace{\sum_{j=1}^{m}\partial_{v}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})K(\tilde{x}_{j},\cdot)}^{\xi}\right\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|a\|_{\mathcal{H}}^{2}.$
@@
Therefore, it is possible to recover $\hat{a}_{\omega}$ by minimizing $a\mapsto L_{adj}(\omega,a)$ over the span of $(\xi,K(x_{1},\cdot),\ldots,K(x_{n},\cdot))$. Hence, to find the optimal coefficients $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\coloneqq((\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega})_{i})_{1\leq i\leq n}$ and $\hat{\beta}_{\omega}$ we first need to express the objective $L_{adj}$ in terms of the coefficients $\operatorname{\boldsymbol{\alpha}}\in\mathbb{R}^{n}$ and $\beta\in\mathbb{R}$ for a given $a^{\operatorname{\boldsymbol{\alpha}},\beta}\in\mathcal{H}$ of the form $a^{\operatorname{\boldsymbol{\alpha}},\beta}=\sum_{i=1}^{n}(\operatorname{\boldsymbol{\alpha}})_{i}K(x_{i},\cdot)+\beta\xi$. To this end, note that the vector $(\xi(x_{1}),\ldots,\xi(x_{n}))$ is exactly equal to $\mathbf{u}=\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}}$ as defined in Proposition B.1. Moreover, using the reproducing property, we directly have:
$(a^{\operatorname{\boldsymbol{\alpha}},\beta}(x_{i}))_{1\leq i\leq n}=\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix},\qquad\langle a^{\operatorname{\boldsymbol{\alpha}},\beta},\xi\rangle_{\mathcal{H}}=\begin{bmatrix}\mathbf{u}^{\top}&{\left\lVert\xi\right\rVert}^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix},\qquad{\left\lVert a^{\operatorname{\boldsymbol{\alpha}},\beta}\right\rVert}^{2}_{\mathcal{H}}=\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\\ \mathbf{u}^{\top}&{\left\lVert\xi\right\rVert}^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}.$
@@
$\widehat{L}_{adj}(\omega,a^{\operatorname{\boldsymbol{\alpha}},\beta})=\frac{1}{2n}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}\\ \mathbf{u}^{\top}\end{bmatrix}\operatorname{\mathbf{D_{v,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}+\frac{1}{m}\begin{bmatrix}\mathbf{u}^{\top}&{\left\lVert\xi\right\rVert}^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}+\frac{\lambda}{2}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\\ \mathbf{u}^{\top}&{\left\lVert\xi\right\rVert}^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}.$
@@
Expression of ^$\widehat{\nabla\mathcal{F}}(\omega)$. The result follows directly after expressing ^$\widehat{\nabla\mathcal{F}}(\omega)$ in vector form using the notations $\operatorname{\mathbf{D_{\omega}^{out}}}$ and $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ from Proposition B.1 and recalling that $(\hat{a}_{\omega}(x_{i}))_{1\leq i\leq n}=\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}$. ∎
Proof of Proposition 3.1.
@@
^$\widehat{\nabla\mathcal{F}}(\omega)=$ $\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}),$
where $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ are given by Equations 10 and 13. We will show that ^$\widehat{\nabla\mathcal{F}}(\omega)=\nabla\widehat{\mathcal{F}}(\omega)$. By Lemmas B.2 and B.1, we know that ^$\nabla\widehat{\mathcal{F}}(\omega)$ and ^$\widehat{\nabla\mathcal{F}}(\omega)$ admit the following expressions:
^$\nabla\widehat{\mathcal{F}}(\omega)$ $=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}-\frac{1}{m}\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}\mathbf{u}$
^$\widehat{\nabla\mathcal{F}}(\omega)$ $=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{n}\operatorname{\mathbf{D_{\omega,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}.$
@@
^$\widehat{\nabla\mathcal{F}}(\omega)-\nabla\widehat{\mathcal{F}}(\omega)$ $=\frac{1}{m}\operatorname{\mathbf{D_{\omega,v}^{in}}}{\left(\operatorname{\mathbf{M}}^{-1}\mathbf{u}+\frac{m}{n}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}\right)}$
@@
where the term $\mathbf{u}+\frac{m}{n}(\operatorname{\mathbf{M}}\operatorname{\mathbf{K}}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}+\hat{\beta}_{\omega}\operatorname{\mathbf{M}}\mathbf{u})$ is equal to 00 by definition of $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}$ and $\hat{\beta}_{\omega}$ as solutions of the linear system (12) of Lemma B.2. ∎
Appendix C Preliminary Results
In this section, $\Omega$ is an arbitrary compact subset of $\mathbb{R}^{d}$ with $\text{hull}(\Omega)$ denoting its convex hull, which is also compact. We also consider an arbitrary fixed positive value $\Lambda$ such that $\lambda\leq\Lambda$ as this would allow us to simplify the dependence of the boundedness and Lipschitz constants on $\lambda$.
C.1 Boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$
Proposition C.1 (Boundedness of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$).
Under Assumptions (B), (D), (A) and (C), the functions $\omega\mapsto\|h^{\star}_{\omega}\|_{\mathcal{H}}$ and $\omega\mapsto\|\hat{h}_{\omega}\|_{\mathcal{H}}$ are bounded over $\textnormal{hull}(\Omega)$ by $\frac{B\sqrt{\kappa}}{\lambda}$, where $B\coloneqq\sup_{\omega\in\textnormal{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. Moreover, for all $\omega\in\textnormal{hull}(\Omega)$ and $x\in\mathcal{X}$, $h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ take value in the compact interval $\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$.

### v3

@@ Appendix C Gradient Estimators
Let $\omega\in\mathbb{R}^{d}$. Recall the expression of $\widehat{\mathcal{F}}(\omega)$:
$\widehat{\mathcal{F}}(\omega)\coloneqq\frac{1}{m}\sum_{j=1}^{m}\ell_{out}\left(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j}\right)$
@@
$\hat{h}_{\omega}=\sum_{i=1}^{n}(\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega})_{i}K(x_{i},\cdot).$ (9)
Hence, finding $\hat{h}_{\omega}$ amounts to minimizing $\widehat{L}_{in}(\omega,h)$ over the span of $(K(x_{1},\cdot),\ldots,K(x_{n},\cdot))$, i.e., over functions $h^{\operatorname{\boldsymbol{\gamma}}}$ of the form $h^{\operatorname{\boldsymbol{\gamma}}}=\sum_{i=1}^{n}(\operatorname{\boldsymbol{\gamma}})_{i}K(x_{i},\cdot)$ for $\operatorname{\boldsymbol{\gamma}}\in\mathbb{R}^{n}$. Restricting the objective to such functions results in the following inner optimization problem which is finite-dimensional:
$\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}\coloneqq\operatorname*{arg\,min}_{\operatorname{\boldsymbol{\gamma}}\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell_{in}\left(\omega,(\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}})_{i},y_{i}\right)+\frac{\lambda}{2}\operatorname{\boldsymbol{\gamma}}^{\top}\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}},$
where we used that $(h^{\operatorname{\boldsymbol{\gamma}}}(x_{i}))_{1\leq i\leq n}=\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}}$ and $\left\lVert h^{\operatorname{\boldsymbol{\gamma}}}\right\rVert^{2}_{\mathcal{H}}=\operatorname{\boldsymbol{\gamma}}^{\top}\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}}$. Similarly, using that $(h^{\operatorname{\boldsymbol{\gamma}}}(\tilde{x}_{j}))_{1\leq j\leq m}=\operatorname{\mathbf{\overline{K}}}\operatorname{\boldsymbol{\gamma}}$, we can express $\widehat{\mathcal{F}}(\omega)$ as follows:
$\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\sum_{j=1}^{m}\ell_{out}\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega})_{j},\tilde{y}_{j}\right).$
@@
$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega})_{j},\tilde{y}_{j}\right)+\frac{1}{m}\sum_{j=1}^{m}(\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}\operatorname{\mathbf{\overline{K}}}^{\top})_{j}\partial_{v}\ell_{out}\left(\omega,(\operatorname{\mathbf{\overline{K}}}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega})_{j},\tilde{y}_{j}\right),$
where $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ denotes the Jacobian of $\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$. We can further express the above equation in matrix form to get:
$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{m}\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}}.$ (10)
Moreover, an application of the implicit function theorem [^1] In the case where the matrix $\operatorname{\mathbf{K}}$ is non-invertible, one needs to restrict $\operatorname{\boldsymbol{\gamma}}$ to the orthogonal complement of the null space of $\operatorname{\mathbf{K}}$. Such a restriction is valid since the resulting solution $\hat{h}_{\omega}$ will not depend on the component belonging to the null space of $\operatorname{\mathbf{K}}$. allows to directly express the Jacobian $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ as a solution of the following linear system obtained by differentiating the optimality condition for $\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ w.r.t. $\omega$:
@@
A solution of the form $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}=-\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}$ always exists by invertibility of the matrix $\operatorname{\mathbf{M}}$. The result follows after replacing $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ by $\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}$ in Equation˜10. ∎
Lemma C.2 (Estimator of the total functional gradient).
@@
$\widehat{\nabla\mathcal{F}}(\omega)=$ $\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}).$
Then, under Assumptions˜(A), (C), (E), (B) and (D), $\widehat{\nabla\mathcal{F}}(\omega)$ admits the following expression:
$\widehat{\nabla\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{n}\operatorname{\mathbf{D_{\omega,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}\in\mathbb{R}^{d},$
where $\operatorname{\mathbf{D_{v}^{out}}},\operatorname{\mathbf{D_{v,v}^{in}}},\operatorname{\mathbf{D_{\omega}^{out}}}$, and $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ are the same matrices given in Proposition˜C.1, while $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\in\mathbb{R}^{n}$ and $\hat{\beta}_{\omega}\in\mathbb{R}$ are solutions to the linear system:
@@
where the vector $\mathbf{u}$ and matrix $\operatorname{\mathbf{M}}$ are the same as in Proposition˜C.1, while $p$ and $v$ are non-negative scalars.
@@
Let $\omega\in\mathbb{R}^{d}$. We start by providing an expression of $\hat{a}_{\omega}$ as a linear combination of the kernel evaluated at the inner training points $x_{i}$, i.e., $K(x_{i},\cdot)$, and some element $\xi\in\mathcal{H}$ that we will characterize shortly. From it, we will obtain the expression of $\widehat{\nabla\mathcal{F}}(\omega)$.
Expression of $\hat{a}_{\omega}$. Recall that $\hat{a}_{\omega}$ is the unique minimizer of $\widehat{L}_{adj}$ in Equation˜5, which admits, for any $a\in\mathcal{H}$, the following simple expression by the reproducing property:
$\widehat{L}_{adj}(\omega,a)=\frac{1}{2n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})a^{2}(x_{i})\\ +\frac{1}{m}\left\langle a,\overbrace{\sum_{j=1}^{m}\partial_{v}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})K(\tilde{x}_{j},\cdot)}^{\xi}\right\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|a\|_{\mathcal{H}}^{2}.$
@@
Therefore, it is possible to recover $\hat{a}_{\omega}$ by minimizing $a\mapsto L_{adj}(\omega,a)$ over the span of $(\xi,K(x_{1},\cdot),\ldots,K(x_{n},\cdot))$. Hence, to find the optimal coefficients $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\coloneqq((\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega})_{i})_{1\leq i\leq n}$ and $\hat{\beta}_{\omega}$, we first need to express the objective $L_{adj}$ in terms of the coefficients $\operatorname{\boldsymbol{\alpha}}\in\mathbb{R}^{n}$ and $\beta\in\mathbb{R}$ for a given $a^{\operatorname{\boldsymbol{\alpha}},\beta}\in\mathcal{H}$ of the form $a^{\operatorname{\boldsymbol{\alpha}},\beta}=\sum_{i=1}^{n}(\operatorname{\boldsymbol{\alpha}})_{i}K(x_{i},\cdot)+\beta\xi$. To this end, note that the vector $(\xi(x_{1}),\ldots,\xi(x_{n}))$ is exactly equal to $\mathbf{u}=\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}}$ as defined in Proposition˜C.1. Moreover, using the reproducing property, we directly have:
$(a^{\operatorname{\boldsymbol{\alpha}},\beta}(x_{i}))_{1\leq i\leq n}=\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix},\langle a^{\operatorname{\boldsymbol{\alpha}},\beta},\xi\rangle_{\mathcal{H}}=\begin{bmatrix}\mathbf{u}^{\top}&\left\lVert\xi\right\rVert^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}$,
$\left\lVert a^{\operatorname{\boldsymbol{\alpha}},\beta}\right\rVert^{2}_{\mathcal{H}}=\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\\ \mathbf{u}^{\top}&\left\lVert\xi\right\rVert^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}$.
@@
$\widehat{L}_{adj}(\omega,a^{\operatorname{\boldsymbol{\alpha}},\beta})=\frac{1}{2n}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}\\ \mathbf{u}^{\top}\end{bmatrix}\operatorname{\mathbf{D_{v,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}\\ +\frac{1}{m}\begin{bmatrix}\mathbf{u}^{\top}&\left\lVert\xi\right\rVert^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}+\frac{\lambda}{2}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}^{\top}&\beta\end{bmatrix}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\\ \mathbf{u}^{\top}&\left\lVert\xi\right\rVert^{2}_{\mathcal{H}}\end{bmatrix}\begin{bmatrix}\operatorname{\boldsymbol{\alpha}}\\ \beta\end{bmatrix}$.
@@
Expression of $\widehat{\nabla\mathcal{F}}(\omega)$. The result follows directly after expressing $\widehat{\nabla\mathcal{F}}(\omega)$ in vector form using the notations $\operatorname{\mathbf{D_{\omega}^{out}}}$ and $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ from Proposition˜C.1 and recalling that $(\hat{a}_{\omega}(x_{i}))_{1\leq i\leq n}=\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}$. ∎
Proof of Proposition˜3.1.
@@
$\widehat{\nabla\mathcal{F}}(\omega)=$ $\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}),$
where $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ are given by Equations˜9 and 12. We will show that $\widehat{\nabla\mathcal{F}}(\omega)=\nabla\widehat{\mathcal{F}}(\omega)$. By Propositions˜C.1 and C.2, we know that $\nabla\widehat{\mathcal{F}}(\omega)$ and $\widehat{\nabla\mathcal{F}}(\omega)$ admit the following expressions:
$\nabla\widehat{\mathcal{F}}(\omega)$ $=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}-\frac{1}{m}\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}\mathbf{u}$
$\widehat{\nabla\mathcal{F}}(\omega)$ $=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}+\frac{1}{n}\operatorname{\mathbf{D_{\omega,v}^{in}}}\begin{bmatrix}\operatorname{\mathbf{K}}&\mathbf{u}\end{bmatrix}\begin{bmatrix}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}\\ \hat{\beta}_{\omega}\end{bmatrix}$.
@@
[…]
@@
where the term $\mathbf{u}+\frac{m}{n}(\operatorname{\mathbf{M}}\operatorname{\mathbf{K}}\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}+\hat{\beta}_{\omega}\operatorname{\mathbf{M}}\mathbf{u})$ is equal to 0 by definition of $\hat{\operatorname{\boldsymbol{\alpha}}}_{\omega}$ and $\hat{\beta}_{\omega}$ as solutions of the linear system (11) of Lemma˜C.2. ∎
Appendix D Preliminary Results
In this section, $\Omega$ is an arbitrary compact subset of $\mathbb{R}^{d}$ with $\text{hull}(\Omega)$ denoting its convex hull, which is also compact. We also consider an arbitrary fixed positive value $\Lambda$ such that $\lambda\leq\Lambda$ as this would allow us to simplify the dependence on $\lambda$ of the boundedness and Lipschitz constants.
D.1 Boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$
Proposition D.1 (Boundedness of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$).
Under Assumptions˜(A), (C), (E), (B) and (D), the functions $\omega\mapsto\|h^{\star}_{\omega}\|_{\mathcal{H}}$ and $\omega\mapsto\|\hat{h}_{\omega}\|_{\mathcal{H}}$ are bounded over $\textnormal{hull}(\Omega)$ by $\frac{B\sqrt{\kappa}}{\lambda}$, where $B\coloneqq\sup_{\omega\in\textnormal{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. Moreover, for all $\omega\in\textnormal{hull}(\Omega)$ and $x\in\mathcal{X}$, $h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ take value in the compact interval $\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$.